# An Online Learning Approach to Prompt-based Selection of Generative Models and LLMs

### v3

@@ 6.1 Summary of the Numerical Results
The results of our numerical experiments indicate the improvement of the proposed PAK-UCB algorithm over the one-arm oracle baseline that is aware of the single generative model with the maximized averaged scores over the prompt distribution. This result suggests that the online learning algorithm could outperform a selector with the side-knowledge of the single best-performing model. We note that this improvement follows the prompt-based selection design of our proposed PAK-UCB algorithm. Moreover, our numerical results indicate that PAK-UCB can significantly improve upon the online learning baselines with shared weights, including LinUCB (Chu et al.,, 2011) and KernelUCB (Valko et al.,, 2013) methods, indicating the effectiveness of the arm-specific design of PAK-UCB in the prompt-based model selection setting. Finally, in our experiments, the proposed RFF-UCB-RBF variant could reduce the computational costs of the general PAK-UCB algorithm.
@@ 6.2 Experiment Settings
Setup 2. Prompt-based selection of LLMs. In these experiments, we focus on prompt-based selection of large language models (LLMs) on various tasks, including Sudoku-solving and code generation (code completion and translation). In the first setup, we select LLMs to solve code generation tasks. The model is given a Python code completion problem (sampled from the first $164$ tasks in BigCodeBench (Zhuo et al.,, 2024)) or C++-to-Java code translation problem (sampled from the HumanEval-X benchmark (Zheng et al.,, 2023)). Sample prompts can be found in Figure 13 and 14 in the Appendix. Particularly, the Gemini-2.5-Flash-preview model attains a higher pass@1 rate on code completion ($55.91$% versus $13.84$%) while underperforms for C++-to-Java translation ($24.39$% versus $40.91$%) compared to the Qwen-Plus model. The rewards are binary, indicating whether the generated code can pass all the test cases. The input of the PAK-UCB, Naive-KRR, and RFF-UCB-RBF methods is the $768$-dimensional embedded prompt that is output by the RoBERTa model (for code completion) or the CodeBERT model [^8] https://huggingface.co/microsoft/codebert-base (for code translation). For LinUCB and KernelUCB baselines, we also concatenate the one-hot encoded vector of the model index to the embedded prompt. The results (Figures 4(a) and LABEL:fig:code-gen-full) show that the proposed RFF-UCB-RBF method queries from the better model conditioned to the task category (Figure LABEL:fig:llm-selection).
@@ 7 Conclusion
In this work, we investigated prompt-based selection of generative models using a contextual bandit algorithm, which can identify the best available generative model for a given text prompt. We propose two kernel-based algorithms, including PAK-UCB and RFF-UCB, to perform this selection task. Our numerical results on LLMs (text-to-text), text-to-image, text-to-video, and image-captioning tasks demonstrate the effectiveness of the proposed framework in scenarios where the available generative models have varying performance rankings depending on the type of prompt. A potential future direction is to extend the online selection to capture the diversity and novelty of generated data. Also, developing cost-aware online selection methods, similar to the cost-aware CB algorithm in (Hu et al., 2025b, ), will be a relevant direction for future studies.
@@
The technical challenge in analyzing PAK-UCB is that predictions in later iterations make use of previous outcomes. Hence, the rewards $\{s_{i}\}_{i\in\Psi_{g}}$ are not independent if the index set $\Phi_{g}$ is updated each time when model $g$ is chosen (Line 6 of Algorithm 2). To address this problem, we leverage a standard approach used in prior works (Auer,, 2003; Chu et al.,, 2011; Valko et al.,, 2013) and present a variant of PAK-UCB in Algorithm 3, which is called Sup-PAK-UCB.
@@ A.1 Regret Analysis of Sup-PAK-UCB
from Line 13 of Algorithm 3. In addition, by a similar statement of (Valko et al.,, 2013, Lemma 4), which is stated in Lemma 5, we have that
@@ A.2 Auxiliary Lemmas
Lemma 2 (Auer, (2003), Lemma 14).
@@
Lemma 5 (Valko et al., (2013), Lemma 4).
@@ B.3 Auxiliary Lemmas
1. Bounding $|(\widetilde{k}_{y}-k_{y})^{\top}Qv|$. Note that ${\mathcal{Y}}\subset{\mathbb{S}}^{d-1}$. We evoke (Sutherland and Schneider,, 2015, Proposition 1), which is rewritten in Lemma 11 using our notations. For a desired threshold $\operatorname{\epsilon_{\textup{RFF}}}>0$, set
@@
where the first inequality holds by the fact that $(\varphi(y))^{\top}\varphi(y_{i})\leq 1$. To bound $\|\widetilde{Q}-Q\|_{2}$, we evoke (Avron et al.,, 2017, Theorem 7), which is rewritten in Lemma 10. For a desired threshold $\operatorname{\Delta_{\textup{RFF}}}\leq 1/2$, the following inequality holds with probability at least $1-\frac{\delta}{2}$:
@@
Lemma 10 (Avron et al., (2017), Theorem 7).
@@
Lemma 11 (Adapted from (Sutherland and Schneider,, 2015, Proposition 1)).
@@ C.1 Results on Setups 1 and 2
Figure 7: Prompt-based generated images from UniDiffuser (Bao et al., 2023a, ) and PixArt-$\alpha$: UniDiffuser attains a higher CLIPScore in generating type “train” prompts (35.29 versus 34.25) while underperforms for type “baseball bat” prompts (32.51 versus 34.30).
@@ C.3 Results on Other Conditional Generation Tasks
3. Text-to-Video (T2V). We provide numerical results on a synthetic T2V setting. Specifically, both the captions and videos are randomly selected from the following five categories of the MSR-VTT dataset (Xu et al.,, 2016): sports/action, movie/comedy, vehicles/autos, music, and food/drink. Each of the five synthetic arms corresponds to an expert in “generating” videos from a single category. Gaussian noises are applied to the video for non-experts. The results are summarized in Figure LABEL:fig:t2v.

### v4

@@ 6.1 Summary of the Numerical Results
The results of our numerical experiments indicate the improvement of the proposed PAK-UCB algorithm over the one-arm oracle baseline that is aware of the single generative model with the maximized averaged scores over the prompt distribution. This result suggests that the online learning algorithm could outperform a selector with the side-knowledge of the single best-performing model. We note that this improvement follows the prompt-based selection design of our proposed PAK-UCB algorithm. Moreover, our numerical results indicate that PAK-UCB can significantly improve upon the online learning baselines with shared weights, including LinUCB (Chu et al., 2011) and KernelUCB (Valko et al., 2013) methods, indicating the effectiveness of the arm-specific design of PAK-UCB in the prompt-based model selection setting. Finally, in our experiments, the proposed RFF-UCB-RBF variant could reduce the computational costs of the general PAK-UCB algorithm.
@@ 6.2 Experiment Settings
Setup 2. Prompt-based selection of LLMs. In these experiments, we focus on prompt-based selection of large language models (LLMs) on various tasks, including Sudoku-solving and code generation (code completion and translation). In the first setup, we select LLMs to solve code generation tasks. The model is given a Python code completion problem (sampled from the first $164$ tasks in BigCodeBench (Zhuo et al., 2024)) or C++-to-Java code translation problem (sampled from the HumanEval-X benchmark (Zheng et al., 2023)). Sample prompts can be found in Figure 13 and 14 in the Appendix. Particularly, the Gemini-2.5-Flash-preview model attains a higher pass@1 rate on code completion ($55.91$% versus $13.84$%) while underperforms for C++-to-Java translation ($24.39$% versus $40.91$%) compared to the Qwen-Plus model. The rewards are binary, indicating whether the generated code can pass all the test cases. The input of the PAK-UCB, Naive-KRR, and RFF-UCB-RBF methods is the $768$-dimensional embedded prompt that is output by the RoBERTa model (for code completion) or the CodeBERT model [^8] https://huggingface.co/microsoft/codebert-base (for code translation). For LinUCB and KernelUCB baselines, we also concatenate the one-hot encoded vector of the model index to the embedded prompt. The results (Figures 4(a) and LABEL:fig:code-gen-full) show that the proposed RFF-UCB-RBF method queries from the better model conditioned to the task category (Figure LABEL:fig:llm-selection).
@@ 7 Conclusion
In this work, we investigated prompt-based selection of generative models using a contextual bandit algorithm, which can identify the best available generative model for a given text prompt. We propose two kernel-based algorithms, including PAK-UCB and RFF-UCB, to perform this selection task. Our numerical results on LLMs (text-to-text), text-to-image, text-to-video, and image-captioning tasks demonstrate the effectiveness of the proposed framework in scenarios where the available generative models have varying performance rankings depending on the type of prompt. A potential future direction is to extend the online selection to capture the diversity and novelty of generated data. Also, developing cost-aware online selection methods, similar to the cost-aware CB algorithm in (Hu et al., 2025b), will be a relevant direction for future studies.
Acknowledgements
The work of Farzan Farnia is partially supported by a grant from the Research Grants Council of the Hong Kong Special Administrative Region, China, Project 14209920, and is partially supported by CUHK Direct Research Grants with CUHK Project No. 4055164 and 4937054. Also, the authors would like to thank the anonymous reviewers for their constructive feedback and suggestions.
Impact Statement
This paper presents work whose goal is to advance the field of Machine Learning. There are many potential societal consequences of our work, none which we feel must be specifically highlighted here.
@@ Acknowledgements
The technical challenge in analyzing PAK-UCB is that predictions in later iterations make use of previous outcomes. Hence, the rewards $\{s_{i}\}_{i\in\Psi_{g}}$ are not independent if the index set $\Phi_{g}$ is updated each time when model $g$ is chosen (Line 6 of Algorithm 2). To address this problem, we leverage a standard approach used in prior works (Auer, 2003; Chu et al., 2011; Valko et al., 2013) and present a variant of PAK-UCB in Algorithm 3, which is called Sup-PAK-UCB.
@@ A.1 Regret Analysis of Sup-PAK-UCB
from Line 13 of Algorithm 3. In addition, by a similar statement of (Valko et al., 2013, Lemma 4), which is stated in Lemma 5, we have that
@@ A.2 Auxiliary Lemmas
Lemma 2 (Auer (2003), Lemma 14).
@@
Lemma 5 (Valko et al. (2013), Lemma 4).
@@ B.3 Auxiliary Lemmas
1. Bounding $|(\widetilde{k}_{y}-k_{y})^{\top}Qv|$. Note that ${\mathcal{Y}}\subset{\mathbb{S}}^{d-1}$. We evoke (Sutherland & Schneider, 2015, Proposition 1), which is rewritten in Lemma 11 using our notations. For a desired threshold $\operatorname{\epsilon_{\textup{RFF}}}>0$, set
@@
where the first inequality holds by the fact that $(\varphi(y))^{\top}\varphi(y_{i})\leq 1$. To bound $\|\widetilde{Q}-Q\|_{2}$, we evoke (Avron et al., 2017, Theorem 7), which is rewritten in Lemma 10. For a desired threshold $\operatorname{\Delta_{\textup{RFF}}}\leq 1/2$, the following inequality holds with probability at least $1-\frac{\delta}{2}$:
@@
Lemma 10 (Avron et al. (2017), Theorem 7).
@@
Lemma 11 (Adapted from (Sutherland & Schneider, 2015, Proposition 1)).
@@ C.1 Results on Setups 1 and 2
Figure 7: Prompt-based generated images from UniDiffuser (Bao et al., 2023a) and PixArt-$\alpha$: UniDiffuser attains a higher CLIPScore in generating type “train” prompts (35.29 versus 34.25) while underperforms for type “baseball bat” prompts (32.51 versus 34.30).
@@ C.3 Results on Other Conditional Generation Tasks
3. Text-to-Video (T2V). We provide numerical results on a synthetic T2V setting. Specifically, both the captions and videos are randomly selected from the following five categories of the MSR-VTT dataset (Xu et al., 2016): sports/action, movie/comedy, vehicles/autos, music, and food/drink. Each of the five synthetic arms corresponds to an expert in “generating” videos from a single category. Gaussian noises are applied to the video for non-experts. The results are summarized in Figure LABEL:fig:t2v.